# Cross-Subject DD: A Cross-Subject Brain-Computer Interface Algorithm

### v1

@@ 2.2.1 Symbolic simplification of the DD model
To generate individualized relation spectrum for multi-subject analysis, the linear classification module is modified by replacing the class-specific weight matrix $W_{2}\in\mathbb{R}^{16\times O}$ with a subject-specific vector $W_{3}\in\mathbb{R}^{16\times 1}$, where all elements are set to 1. This operation consolidates the nonlinear interaction items across hidden units, enabling explicit extraction of per-subject relation spectrum. The revised model is formalized as:
@@
$\left\{\begin{array}[]{l}G_{1}=W_{1}^{T}X\\[8.0pt] G_{2}=W_{G_{1}}G_{1}\circ G_{1}\\[8.0pt] \widetilde{G}^{\text{SSTL-PFoutputs}}=W_{3}^{T}\left(W_{G_{1}}G_{1}\circ G_{1}\right)\end{array}\right.$ (4)
@@
where $\widetilde{G}^{SSTL-PFoutputs}\in\mathbb{R}^{1\times 1}$ represents the revised outputs of SSTL-PF, which are used for transforming relation spectrums. And the expanded output can also be expressed as follows. The relation spectrum is derived by symbolically expanding equation.
@@
$\left\{\begin{array}[]{l}\alpha_{i}=\sum_{k=1}^{16}W_{G_{1}}^{(k,k)}(W_{1}^{(k,i)})^{2}\\[8.0pt] \beta_{ij}=\sum_{k=1}^{16}\sum_{l\neq k}W_{G_{1}}^{(k,l)}W_{1}^{(k,i)}W_{1}^{(l,j)}\\[8.0pt] \widetilde{G}^{\text{SSTL-PFoutputs}}=\sum_{i=1}^{32}\alpha_{i}x_{i}^{2}+\sum_{1\leq i<j\leq 32}\beta_{ij}x_{i}x_{j}\end{array}\right.$ (5)
This polynomial is automated using symbolic tools (e.g., MATLAB Symbolic Toolbox).
2.2.2 Controlling the number of common features
@@
The completeness of the relation spectrum is mathematically guaranteed through its construction, which explicitly preserves all quadratic interactions in the feature space. For a d-dimensional input vector $X\in\mathbb{R}^{d\times 1}$, the relation spectrum comprises $d$ squared items ($x_{i}^{2}$) and $\frac{d(d-1)}{2}$ pairwise interaction items ($x_{i}x_{j}$), resulting in a deterministic theoretical term count.
@@
This constraint is enforced during symbolic simplification, where the dendrite network’s hierarchical operations are systematically expanded into polynomial form. Redundant items are eliminated through algebraic consolidation, while coefficients of identical items are merged. Consequently, the final spectrum contains precisely $N_{t}$ non-redundant items, ensuring exhaustive representation of nonlinear feature interactions without information loss.
2.3 Extraction of common features based on statistical analysis (ECF-SA)
Statistical analysis is employed to identify significant patterns and relationships within data. In this paper, it is employed to extract common features from $N_{\mathrm{s}}-1$ relation spectrums corresponding to $N_{\mathrm{s}}-1$ subjects, allowing for the identification of common components across subjects.
For each quadratic term $\alpha_{i}x_{i}^{2}$ or $\beta_{ij}x_{i}x_{j}$ interaction term in the spectrum, coefficient values $\left\{\alpha_{i}^{(1)},\alpha_{i}^{(2)},\ldots,\alpha_{i}^{(N_{s}-1)}\right\}$ or $\left\{\beta_{ij}^{(1)},\beta_{ij}^{(2)},\ldots,\beta_{ij}^{(N_{s}-1)}\right\}$ are collected across subjects. A one-sample $t$-test is applied to determine whether the mean coefficient significantly deviates from zero. The null and alternative hypotheses are defined as follows.
@@ 2.3 Extraction of common features based on statistical analysis (ECF-SA)
The t-statistic is calculated as:
@@
where $\overline{R_{i}}$ and $s_{i}$ are the sample mean and standard deviation of the coefficient values, $N_{s}-1$ is the number of subjects. Relation items with p-values below a predefined threshold $\alpha_{\text{threshold}}$ are retained as statistically significant common features.
As summarized in Table 1 , common features are characterized by consistent patterns across subjects, whereas subject-specific features exhibit more random and individual-specific variation.In this context, the features correspond to the relation items.
Table 1: Definitions of common features and subject-specific features
@@
common features Features that are consistent across subjects, where the relationships are similar (mostly positive or negative). The accumulated average significantly differs from zero, and the random component is non-zero.
subject-specific features Features that are specific to each subject, with relationships that are random (can be positive or negative). The accumulated average approaches zero, and the random component is zero.
2.4 Building cross-subject decoding model based on common features (BCSDM-CF)
Using the extracted common features across subjects, a generalized BCI model is then constructed.
2.4.1 Feature transformation layer
The common features are derived by transforming each subject’s signal space. To map all subjects’ information into a unified feature space, a feature transformation layer is designed. This layer transforms the input brain signal $X\in\mathbb{R}^{d\times 1}$ into a new feature vector $Z\in\mathbb{R}^{m\times 1}$ , where $m\leq N_{t}$ and $N_{t}=d+\frac{d(d-1)}{2}$ is the theoretical maximum relation items. The transformation involves two types of operations: squared items and interaction items. For each significant squared item $x_{i}^{2}$ , it can be described as follows.
@@ 2.4.1 Feature transformation layer
where the pair $(i,j)$ is included in the set $S_{\text{int}}$ of significant interactions. The transformed feature vector $Z$ is constructed by concatenating all retained components.
@@
where $m=m_{\text{sq}}+m_{\text{int}}$ , ensuring that only cross-subject common features are preserved.
2.4.2 Linear classifier
The linear classifier then maps the m-dimensional feature vector $Z$ to class probabilities through a linear transformation followed by softmax activation. The linear transformation is defined as follows.
$h=\mathbf{W}^{\top}\mathbf{Z}+b$ (12)
where $\mathbf{W}\in\mathbb{R}^{m\times O}$ is the weight matrix, $\mathbf{b}\in\mathbb{R}^{O\times 1}$ is the bias vector, and $O$ is the number of categories. The output probabilities are computed via the softmax function. The model is trained by minimizing the cross-entropy loss.
@@ 3.1 Experimental Details
3.1.1 Dataset
To assess the feasibility and generalizability of CSDD, it is evaluated using a publicly available dataset. The BCIC IV 2a dataset contains EEG signals from 22 electrodes, recorded from nine healthy subjects over two sessions conducted on two different days. Each session consists of 288 four-second motor imagery trials per subject, involving the imagined movements of the left hand, right hand, feet, and tongue. This study uses data from 0 to 4 seconds. Before release, the signals were sampled at 250 Hz and bandpass filtered between 0.5 Hz and 100 Hz. The original dataset uses 288 trials from the first session (“T”) for training and 288 trials from the second session (“E”) for testing. However, in the cross-subject scenario, the original dataset needs to be re-split by subject using a leave-one-subject-out approach for modeling based on common features. Consequently, nine evaluation datasets (A01-A09) are obtained, each containing 576 trials (288 trials $\times$ 2 sessions) for testing one subject, with the remaining 4608 trials (288 trials $\times$ 2 sessions $\times$ 8 subjects) used for training. The BCIC IV 2a dataset is accessible at http://bnci-horizon-2020.eu/database/data-sets. The experimental paradigm of the BCIC IV 2a dataset is shown in Fig. 3.
Figure 3: The experimental paradigm for BCIC IV 2a
@@ 3.1.2 Performance metric
We evaluate the statistical significance between the proposed model and other baseline models using paired $t$-tests and $p$-values. A $p$-value below 0.01 (**) denotes a statistically significant difference, $p$ < 0.05 (*) indicates a significant difference, and $p$ > 0.05 suggests no significant difference.
@@ 3.1.3 Evaluation baselines
We compare the proposed network with six baseline models.
@@
Figure 4: Cross-subject setting.
@@ 3.2 Experimental Setups
For the upcoming experiments, this paper adopts the Leave-One-Subject-Out (LOSO) validation strategy. This method is widely regarded as the most rigorous approach for cross-subject classification, as it completely excludes certain subjects during training and evaluates the model’s performance on unseen subjects[9][30][31][32]. It provides an objective measure of generalization ability. In contrast, other commonly used strategies, such as mix-up-all-subjects and random selection, either merge data from all subjects or partition the data randomly, which may artificially inflate classification performance and fail to reflect real-world adaptability. Given that an ideal BCI system should decode EEG signals from entirely unseen subjects, leave-n-subjects-out remains the most appropriate choice for evaluating cross-subject generalization[33][34][35].
@@
Subject -i Indicates that data from subject A0i is used for testing, while data from the remaining eight subjects is used for training (e.g., leave-one-subject-out setting) (i = 1, 2, …, 9) (see Fig. 4)
@@ 3.2.1 Personalized model training and acquisition
To eliminate inter-subject variability and extract features directly relevant to the motor imagery task, we designed a transfer learning strategy consisting of two stages: pretraining and fine-tuning.
In the pretraining stage, the LOSO strategy was employed, where EEG data from eight subjects were divided into a training set and a validation set in an 8:2 ratio. A five-fold cross-validation was conducted for iterative model optimization and parameter tuning.
In the fine-tuning stage, the pretrained model was adapted to individual subjects using the training data from eight subjects. This module was then independently optimized for each subject, leading to the generation of eight personalized models. Finally, the fine-tuned model was applied to the test set of the left-out subject, and the classification accuracy of the motor imagery task was evaluated to assess the decoding performance of the personalized model. Hyperparameter optimization was performed using grid search.
During the pretraining stage, the learning rate was set to 0.0002, the batch size was 64, and the model was trained for 500 epochs. In the fine-tuning stage, the learning rate remained at 0.0002, but the batch size was increased to 256 to accelerate the optimization process, with training conducted for 200 epochs.
3.2.2 Extraction of common relation items
Using the best parameters obtained during the training phase, the relation spectrums of the eight personalized models were expanded, resulting in eight corresponding relation spectrums. These relation spectrums are expressed in polynomial form, where the relation items are identical across subjects, but the coefficients vary for each subject. To identify common relation items, a one-sample $t$-test was conducted on each of the 528 relation items to assess the difference between their mean values and zero. By setting different significance levels, we selected the statistically significant relation items as common features.
To investigate the impact of the number of common relation items on cross-subject decoding performance, we analyzed the common relation items with predefined significance thresholds set to $\alpha_{\text{threshold}}$ values of 0.5, 0.1, 0.01, and 0.001.
3.2.3 Building cross-subject models with common features
To validate the cross-subject performance of the proposed CSDD model, we employ LOSO for evaluation.
Specifically, in each experiment, eight subjects are selected as the training set, and the data from the remaining subject are used as the test set, ensuring that the model is evaluated on unseen individuals. Through nine iterations of experiment, the cross-subject decoding results based on the CSDD model are calculated for each subject, and the average performance of the model in unseen individuals is statistically analyzed. In this stage, the learning rate was set to 0.0001, the batch size was set to 32, and the number of epochs was set to 200. The decoding results of different subjects were obtained by selecting the common features that yielded the best decoding performance for each subject.
The statistical significance of the differences in decoding performance achieved by different methods was assessed by the paired sample $t$-test for the BCIC IV 2a dataset.
Figure 5: Transform SSTL-PF models. (Test data for subject 9 and train data for the remaining eight subjects). (a) The relation spectrum of subject 1-subject 8. (b) Correlation matrix between eight personalized models.
4 Results
4.1 Relation spectrum of the personalized model
Fig. 5 (a) depicts the relation spectrum of the personalized models for subject 1 to subject 8, with each curve representing the relation spectrum of an individual subject. The x-axis corresponds to 528 relation items, and the y-axis represents the coefficients associated with these relation items. Upon observation, significant differences can be observed between the relation spectrums of different subjects, reflecting the influence of individual differences on EEG signal decoding. This shows that the model training process successfully captured the features of each subject.
Furthermore, Fig. 5 (b) presents the correlation matrix of the eight personalized models. In the matrix, the intensity of the color blocks indicates the similarity of relation spectrums between subjects, with darker blocks representing higher correlation and lighter blocks representing lower correlation. It can be observed that, while there is some similarity between subjects, individual differences remain prominent. In particular, the relation spectrums of certain subjects are highly similar, highlighting the potential for cross-subject feature sharing. This result further corroborates the effectiveness of the CSDD model in extracting common features and underscores the importance of distinguishing between common and subject-specific features for enhancing cross-subject decoding performance.
4.2 Common Features
4.2.1 Visualization of common features
To validate the effectiveness of the proposed cross-subject decoding method, we performed a key visualization analysis, focusing on the extraction of common features. Feature visualization not only provides an intuitive view of the data distribution but also offers valuable support for subsequent classification tasks.
Fig. 6 displays the extracted common features across different subjects, based on data from subject 1 to subject 8 (Subject-9), along with the relation spectrum analysis results. This visualization allows us to observe the common relationships between different subjects intuitively. These features represent stable components shared by multiple subjects, providing visual evidence to confirm whether common features with broad applicability have been successfully extracted. The figure clearly highlights the stability of the extracted common features, suggesting that they are effective and representative in cross-subject decoding tasks.
Figure 6: Visualization of common and subject-specific features. (e.g. Subject-9)
4.2.2 The number of common features corresponding to different predefined thresholds $\alpha_{\text{threshold}}$
As shown in Fig. 7, the number of common relation items decreases with stricter predefined thresholds $\alpha_{\text{threshold}}$. When $\mathit{p}$ < 0.1, the number of features ranges from 361 to 406 across subjects. As the threshold is reduced to 0.05, 0.01, and 0.001, fewer features are retained, with the number of features dropping to as low as 125 for Subject-3 when $\mathit{p}$ < 0.001. This indicates that stricter thresholds lead to fewer but more statistically significant features.
Figure 7: The change in the number of common relation items in the CSDD on the BCIC IV 2a dataset is observed as the predefined threshold $\alpha_{\text{threshold}}$ is gradually increased from 0.001 to 0.1.
4.3 Decoding Performance Evaluation
@@
CSDD 69.10 49.31 80.38 55.03 63.02 54.17 76.04 70.83 65.80 64.85 10.47 0.531 -

### v2

@@ 2.2.3 Expansion of the Personalized Model (EPM)-Expansion
To generate individualized relation spectrums for multi-subject analysis, the linear classification module is modified by replacing the class-specific weight matrix $\boldsymbol{W}_{2}\in\mathbb{R}^{16\times O}$ with a subject-specific vector $\boldsymbol{W}_{3}\in\mathbb{R}^{16\times 1}$, where all elements are set to 1. The revised model is expressed as follows.
@@
$\left\{\begin{array}[]{l}\boldsymbol{G}_{1}=\boldsymbol{W}_{1}^{T}\boldsymbol{X}\\[8.0pt] \boldsymbol{G}_{2}=\boldsymbol{W}_{G_{1}}\boldsymbol{G}_{1}\circ\boldsymbol{G}_{1}\\[8.0pt] \widetilde{G}^{\text{APM\_outputs}}=\boldsymbol{W}_{3}^{T}(\boldsymbol{W}_{G_{1}}\boldsymbol{G}_{1}\circ\boldsymbol{G}_{1})\end{array}\right$. (8)
@@
where $\widetilde{G}^{APMoutputs}\in\mathbb{R}^{1\times 1}$ represents the revised outputs of APM, which are used for transforming relation spectrums.
The expanded output can be expressed as follows.
@@
$\left\{\begin{array}[]{l}\alpha_{i}=\sum_{k=1}^{16}\boldsymbol{W}_{G_{1}}^{(k,k)}\left(\boldsymbol{W}_{1}^{(k,i)}\right)^{2}\\[8.0pt] \beta_{ij}=\sum_{k=1}^{16}\sum_{l\neq k}\boldsymbol{W}_{G_{1}}^{(k,l)}\,\boldsymbol{W}_{1}^{(k,i)}\,\boldsymbol{W}_{1}^{(l,j)}\\[8.0pt] \widetilde{G}^{\text{APM\_outputs}}=\sum_{i=1}^{32}\alpha_{i}x_{i}^{2}+\sum_{1\leq i<j\leq 32}\beta_{ij}x_{i}x_{j}\end{array}\right$. (9)
@@
Controlling the Number of Common Components: To ensure interpretability and computational feasibility, the number of common components must be rigorously controlled. This is achieved by leveraging the inherent completeness property of the relation spectrum. Specifically, its construction explicitly preserves all quadratic interactions in the feature space. For a d-dimensional input vector $\boldsymbol{X}\in\mathbb{R}^{d\times 1}$, the relation spectrum comprises $d$ squared items ($x_{i}^{2}$) and $\frac{d(d-1)}{2}$ pairwise interaction items ($x_{i}x_{j}$), resulting in a deterministic theoretical term count.
@@
This ensures that, for a 32-dimensional feature vector, the relation spectrum contains exactly 528 distinct terms, encompassing all possible squared and pairwise interaction components.
2.2.4 Extraction of Common Components (ECC)—Filtering
Common components are extracted from the relation spectrums of eight subjects through statistical testing. For each quadratic term $\alpha_{i}x_{i}^{2}$ or $\beta_{ij}x_{i}x_{j}$ interaction term in the spectrum, coefficient values $\left\{\alpha_{i}^{(1)},\alpha_{i}^{(2)},\ldots,\alpha_{i}^{(8)}\right\}$ or $\left\{\beta_{ij}^{(1)},\beta_{ij}^{(2)},\ldots,\beta_{ij}^{(8)}\right\}$ are collected across subjects.
A one-sample $t$-test is applied to determine whether the mean coefficient significantly deviates from zero, thereby implementing the filtering operation of the system filter. The definitions of the common and personalized components are summarized in Table 1. The null and alternative hypotheses are defined as follows.
@@ 2.2.4 Extraction of Common Components (ECC)—Filtering
where $R_{i}$ denotes the set of coefficients corresponding to the $i$-th relation item across different subjects.
The test statistic is calculated as follows.
@@
where $\overline{R_{i}}$ is the sample mean, $s_{i}$ is the sample standard deviation, and $N$ is the number of subjects. A significance threshold $\alpha_{\text{threshold}}$ is set according to the experimental requirements. Items with $p$-values below the threshold were identified as common components, which were retained as the target components, whereas items failing to reach significance were classified as personalized components and filtered out as unnecessary.
Table 1: Definitions of common components and personalized components
@@
common components Relation items that are consistent across subjects, where the relationships are similar (mostly positive or negative). The accumulated average significantly differs from zero, and the random component is non-zero.
personalized components Relation items that are specific to each subject, with relationships that are random (can be positive or negative). The accumulated average approaches zero, and the random component is zero.
2.2.5 Construction of a Cross-subject Decoding Model based on Common Components (CCM)—Reconstruction
The retained common components are used to reconstruct a cross-subject decoding model, which was subsequently applied to unseen subjects to evaluate decoding performance. To map all subjects’ information into a unified feature space, a feature transformation layer is designed. This layer transforms the input brain signal $\boldsymbol{X}\in\mathbb{R}^{d\times 1}$ into a new feature vector $\boldsymbol{Z}\in\mathbb{R}^{m\times 1}$ , where $m\leq N_{t}$ and $N_{t}=d+\frac{d(d-1)}{2}$ is the theoretical maximum relation items.
The transformation involves two types of operations: squared items and interaction items. For each significant squared item $x_{i}^{2}$ , it can be described as follows.
@@ 2.2.5 Construction of a Cross-subject Decoding Model based on Common Components (CCM)—Reconstruction
where the pair $(i,j)$ is included in the set $S_{\text{int}}$ of significant interactions. The final transformed feature vector $\boldsymbol{Z}$ is constructed by concatenating all retained components.
@@
where $m=m_{\text{sq}}+m_{\text{int}}$ , ensuring that only cross-subject common features are preserved. The linear classifier then maps the m-dimensional feature vector $\boldsymbol{Z}$ to class probabilities through a linear transformation followed by softmax activation. The linear transformation is defined as follows.
$\boldsymbol{h}=\boldsymbol{W}^{\top}\boldsymbol{Z}+\boldsymbol{b}$ (16)
where $\boldsymbol{W}\in\mathbb{R}^{m\times O}$ is the weight matrix, $\boldsymbol{b}\in\mathbb{R}^{O\times 1}$ is the bias vector, and $O$ is the number of categories. Finally, the output probabilities are computed via the softmax function. The model is trained by minimizing the cross-entropy loss.
@@ 3.1.1 Performance Metric
We evaluate the statistical significance between the proposed model and other baseline models using paired $t$-tests and $p$-values.
@@ 3.1.2 Evaluation Baselines
We compare the proposed CSD-SF with six baseline models.
@@
Figure 5: Cross-subject experimental setting.
@@ 3.2 Experimental Setups
For the upcoming experiments, this paper adopts the Leave-One-Subject-Out (LOSO) validation strategy. This method is widely regarded as the most rigorous approach for cross-subject classification, as it completely excludes certain subjects during training and evaluates the model’s performance on unseen subjects[33, 25, 26, 27]. It provides an objective measure of generalization ability.
@@
Subject -i Indicates that data from subject A0i is used for testing, while data from the remaining eight subjects is used for training (e.g., leave-one-subject-out setting) (i = 1, 2, …, 9) (see Fig. 5).
Figure 6: Transform APM models. (Test data for subject 9 and train data for the remaining eight subjects). (a) The relation spectrum of subject 1-subject 8. (b) Correlation matrix between eight personalized models.
@@ 3.2.1 Personalized Model Training and Acquisition
Hybrid features obtained through HFE were used as input to the DD for training. Using the LOSO strategy, EEG data from eight subjects were split into training and validation sets at an 8:2 ratio. Five-fold cross-validation was used for iterative optimization, with a learning rate of 0.0002, batch size of 64, and 500 training epochs. For personalized modeling, the trained network was further adapted to each subject using their individual training data, resulting in eight subject-specific models. Hyperparameters were optimized via grid search, with the learning rate kept at 0.0002, batch size increased to 256, and 200 epochs of training. Finally, the subject-specific model was applied to the test set of the left-out subject, and classification accuracy was evaluated to assess the decoding performance of the personalized models.
3.2.2 Extraction of Common components
Based on the optimal parameters from model training, the eight subject-specific models were expanded into their corresponding relation spectrums. A one-sample $t$-test was performed on the coefficients of all 528 relation items to assess whether their mean values across subjects significantly differed from zero. Relation items passing the significance test were regarded as common components, with predefined thresholds of $\alpha_{\text{threshold}}=0.5,0.1,0.01,$ and $0.001$ used to analyze their impact on cross-subject decoding performance.
3.2.3 Building Cross-Subject Models with Common Components
To evaluate the proposed CSD-SF framework under cross-subject conditions, the LOSO strategy was adopted. In each iteration, data from eight subjects were used for training, while the remaining subject was held out for testing, ensuring evaluation on unseen individuals. This process was repeated nine times, and the average decoding accuracy across subjects was reported.
Model training in this stage was conducted with a learning rate of 0.0001, a batch size of 32, and 200 epochs. For each subject, the subset of common components yielding the highest decoding accuracy was selected. Finally, statistical comparisons between different methods were performed using paired-sample $t$-tests on the BCIC IV 2a dataset.
Figure 7: Visualization of common and personalized components. (e.g. Subject-9)
@@ 3.2.3 Building Cross-Subject Models with Common Components
CSD-SF 69.10 49.31 80.38 55.03 63.02 54.17 76.04 70.83 65.80 64.85 10.47 0.531 -
*Note: Chance level for the four-class task is 25%; despite inter-subject variability, CSD-SF consistently surpasses this baseline.
4 Results
4.1 Relation Spectrum of the Personalized Model
Fig. 6(a) depicts the relation spectrum of the personalized models for subject 1 to subject 8, with each curve representing the relation spectrum of an individual subject. The x-axis corresponds to 528 relation items, and the y-axis represents the coefficients associated with these relation items. Upon observation, significant differences can be observed between the relation spectrums of different subjects, reflecting the influence of individual differences on EEG signal decoding. This shows that the model training process successfully captured EEG features of each subject. Furthermore, Fig. 6(b) presents the correlation matrix of the eight personalized models. In the matrix, the intensity of the color blocks indicates the similarity of relation spectrums between subjects, with darker blocks representing higher correlation and lighter blocks representing lower correlation. It can be observed that, while there is some similarity between subjects, individual differences remain prominent. In particular, the relation spectrums of certain subjects are highly similar, highlighting the potential for cross-subject feature sharing.
4.2 Common Components
4.2.1 Visualization of Common Components
To validate the ability of the proposed CSD-SF model to extract common components, we performed a key visualization analysis. Fig. 7 displays the extracted common components across different subjects, based on data from subject 1 to subject 8 (Subject-9), along with the relation spectrum analysis results. This visualization allows us to observe the common relationships between different subjects intuitively. These features represent stable components shared by multiple subjects, providing visual evidence to confirm whether common components with broad applicability have been successfully extracted.
4.2.2 The Number of Common Components Corresponding to Different Predefined Thresholds
As shown in Fig. 8, the number of common relation items decreases with stricter predefined thresholds $\alpha_{\text{threshold}}$. When $\mathit{p}<0.1$, the number of features ranges from 361 to 406 across subjects. As the threshold is reduced to 0.05, 0.01, and 0.001, fewer features are retained, with the number of features dropping to as low as 125 for Subject-3 when $\mathit{p}<0.001$. This indicates that stricter thresholds lead to fewer but more statistically significant components.
Figure 8: The change in the number of common relation items in the CSD-SF on the BCIC IV 2a dataset is observed as the predefined threshold $\alpha_{\text{threshold}}$ is gradually increased from 0.001 to 0.1.
4.3 Decoding Performance Evaluation